Find the price of an American call option on a futures if the current spot price is 30, the exercise price is 25, the futures price is 33.70, the risk-free interest rate is 6 percent, the spot asset can go up by 10 percent or down by 8 percent per period and the call expires in two periods, which is also when the futures expires.
Answer #1

Given ey It is a call option where anderlying asset is futures enercise price (X) = 25 spot price (sp) - 30 Risk free interes1.06 - 0.92 1.1 - 0.92 - 0.18 0.999 77.78% te poobability of lower spot proc probability of e sa Higher Sporpoice |- 77.784 2call payoff = sex m-x memoulet me 00 X Exercise price en [asis Diagram 1] payok @ NODE DE 36.3-25 - 11.3 @ Node € = 30-36-25value @ Nodec on Higher of o payoll @ Node c. 276-25 = 2.6 Discounted value of expected payol, 30 Node F&G le (5-36 x 77.78%]